OpenStudy (nermaljean99):

Any help on some Algebra 2? :)

OpenStudy (nermaljean99):

Quadratic Equations

OpenStudy (isaiah.feynman):

Post the actual question.

OpenStudy (nermaljean99):

I will lol hang on

OpenStudy (nermaljean99):

1.) What are the solutions of the quadratic equation?-24 x^2+11x=

OpenStudy (nermaljean99):

hang on it put that wrong... it is x^2+11x=-24

OpenStudy (isaiah.feynman):

Add the 24 to both sides first.

OpenStudy (dumbcow):

you could factor it or use quadratic formula \[x = \frac{-b \pm \sqrt{b^2 -4ac}}{2a}\]

OpenStudy (isaiah.feynman):

Factoring would be fastest here.

OpenStudy (nermaljean99):

I got x^2+35?

OpenStudy (isaiah.feynman):

What did you do to get that?

OpenStudy (nermaljean99):

well i added the 24 to both sides?

OpenStudy (isaiah.feynman):

Oh no..you can 't add 24 to 11x. They are not the same thing.

OpenStudy (nermaljean99):

oh... then how do I do that?

OpenStudy (isaiah.feynman):

You do this..